Find the indicated partial derivative. f(x, y) = y sin-(xy); fy(2, 4)
Answer #1

we have

f(x,y)=ysin^{-1}(xy)

now,

f_y(x,y)=\frac{\partial }{\partial y}ysin^{-1}(xy)

f_y(x,y)=y\frac{\partial }{\partial y}sin^{-1}(xy)+sin^{-1}(xy)\frac{\partial }{\partial y}y

f_y(x,y)=y \cdot \frac{x}{\sqrt{1-x^2y^2}}+sin^{-1}(xy)

f_y(x,y)=\frac{xy}{\sqrt{1-x^2y^2}}+sin^{-1}(xy)

at (x, y) = (2, 1/4),

f_y\left ( 2,\frac{1}{4} \right )=\frac{2 \cdot \frac{1}{4}}{\sqrt{1-2^2\left ( \frac{1}{4} \right )^2}}+sin^{-1}\left ( 2 \cdot \frac{1}{4} \right )

f_y\left ( 2,\frac{1}{4} \right )=\frac{\frac{1}{2}}{\sqrt{1-\frac{1}{4}}}+sin^{-1}\left ( \frac{1}{2} \right )

f_y\left ( 2,\frac{1}{4} \right )=\frac{\frac{1}{2}}{\sqrt{\frac{3}{4}}}+sin^{-1}\left ( \frac{1}{2} \right )

f_y\left ( 2,\frac{1}{4} \right )=\frac{1}{\sqrt{3}}+\frac{\pi}{6}

f_y\left ( 2,\frac{1}{4} \right )=1.10094